OpenStudy (anonymous):

simplify: squareroot of (4x^2 + 4x^4 + 1)

OpenStudy (lgbasallote):

hint: \[4x^2 + 4x^4 + 1 = (2x^2 + 1)^2\]

OpenStudy (anonymous):

4x2+4x4+1=(2x2+1)2 root((2x2+1)2)= (2x2+1)